# Leveraging Label Information for Stealthy Data Stealing in Vertical Federated Learning

### v1

@@ Appendix A DMAVFL_sync algorithm
Algorithm 3 DMAVFL_sync.
@@ Appendix B Datasets and model descriptions
We illustrate all datasets and models information used in the experiments in the following Table 10, where MLP is composed of a linear and an activation, e.g., ReLU layer, CNN layer is composed a convolution process and a linear layer, and Residual block is constructed by 3 CNN layers and a skip connection.
Table 10. Dataset characteristics and models structure
Dataset Credit RT_IOT2022 MNIST CIFAR-10
Feature size
$d_{a},d_{p}$
11, 12 41, 42 392, 392 512, 512
The size of
training set
19,091 89,920 54,546 45,455
The size of
auxiliary set
1,909 8,573 5,454 4,545
The size of
test set
9,000 24,624 10,000 10,000
$f_{e}(\cdot),f_{a}(\cdot),$ and
$f_{p}(\cdot)$ structure
MLP layer
2
MLP layer
2
CNN layer
2
Residual
block *4
$f_{d}(\cdot)$ structure
MLP layer
3
MLP layer
3
CNN layer
3
CNN layer
3
$D(\cdot)$ structure
MLP layer
3
MLP layer
3
CNN layer
2
CNN layer
4

### v2

@@ Appendix A URVFL_sync algorithm
Algorithm 3 URVFL_sync.
@@ Appendix B Datasets and model descriptions
We illustrate all datasets and models information used in the experiments in the following Table X, where MLP is composed of a linear and an activation, e.g., ReLU layer, CNN layer is composed a convolution process and a linear layer, and Residual block is constructed by 3 CNN layers and a skip connection.
Appendix C Visualization of reconstruction in Tiny imagenet
In addition to visualizing the reconstruction performance on CIFAR-10, we compare the reconstruction performance of AGN, FSHA, and URVFL on the Tiny ImageNet dataset in Table XI.
TABLE X: Dataset characteristics and models structure
Dataset Credit RT_IOT2022 MNIST CIFAR-10 Tiny imagenet
Feature sizes of the adversary
and the passive clients $d_{a},d_{p}$
11, 12 41, 42 392, 392 512, 512 2048, 2048
The size of training dataset 19,091 89,920 54,546 45,455 72,728
The size of the auxiliary dataset 1,909 8,573 5,454 4,545 7,272
The size of the test dataset 9,000 24,624 10,000 10,000 20,000
$f_{e}(\cdot),f_{a}(\cdot),$ and $f_{p}(\cdot)$ structure MLP layer×$\times$2 MLP layer×$\times$2 CNN layer×$\times$2 Residual block×$\times$3 Residual block×$\times$3
$f_{d}(\cdot)$ structure MLP layer×$\times$3 MLP layer×$\times$3 CNN layer×$\times$3 CNN layer×$\times$3 CNN layer×$\times$4
$D(\cdot)$ structure MLP layer×$\times$3 MLP layer×$\times$3 CNN layer×$\times$2 CNN layer $\times$3 CNN layer×$\times$3
Learning rate 1e-4 1e-4 1e-4 1e-3 1e-4
URVFL pretraing epochs 30 40 10 20 30
TABLE XI: Visualization of reconstructed features on the Tiny imagenet dataset with and without the detection SG. Except for the original image, the right side of each image is the reconstructed image. The last column is the average reconstruction error.